# boxthm]Proposition boxthm]Example boxthm]Corollary boxthm]Lemma boxthm]Definition

### v2

@@ 4.1 Theoretical analysis
[Closed-form solution] Under 4.1 and 4.2, the Bi-O problem
@@
The proof of 4.2 is deferred to Section A.1.
@@
Under the assumptions of 4.2, the optimal parameters $\theta^{\star}_{\mathrm{U}^{\star}}$ obtained from the lower-level problem with $\mathrm{U}^{\star}=\frac{\lambda}{2}\Sigma^{-1}$ minimize the reverse KL divergence between $\hat{p}_{\theta}$ and $q$:
@@ 4.2 Empirical validation
In this section, we empirically evaluate the theoretical results from Section 4.1. To this end, we generate synthetic data that satisfy 4.1: $\mathcal{D}=\{({\bf x}_{i},{\bf y}_{i})\}_{i=0}^{N}$, where ${\bf x}_{i}\sim\mathcal{U}([-5,5]^{n})$ ($\mathcal{U}$ denotes the uniform distribution), and ${\bf y}_{i}\sim q(.|{\bf x}_{i}):=\mathcal{N}(\Lambda{\bf x}_{i},\Sigma)$ with diagonal covariance matrix $\Sigma=\beta^{2}\mathrm{I}_{n}$ and $\beta>0$. For the model $\hat{p}_{\theta}$, we relax the linearity and homoscedasticity assumptions by considering a neural network that parametrizes a Gaussian distribution, in which both the mean function and the diagonal covariance matrix depend on the input: $\hat{p}_{\theta}(.|{\bf x}_{i}):=\mathcal{N}(\mu_{\theta}({\bf x}_{i}),\sigma^{2}_{\theta}({\bf x}_{i})\mathrm{I}_{n})$. We compare baselines trained with the negative log-likelihood (NLL) and mean squared error (MSE) losses against PG variants, using either a negative squared distance reward $\mathrm{U}=\mathrm{I}_{n}$ or the optimal reward function derived in 4.2 with $\mathrm{U}^{\star}=\frac{\lambda n}{2\operatorname{Tr}(\Sigma)}\mathrm{I}_{n}$.
@@
Figure 2: Learned distributions comparison on a single data point. The PG loss paired with the optimal reward function in 4.2 shows optimal convergence, even when compared with the baseline directly optimizing the NLL.
@@ 5.2 Empirical validation
Fig. 3 presents the results of running an implicit differentiation solver for 100 outer iterations, each with 50 inner iterations, and a learning rate of $10^{-2}$ for both optimization loops. The outer optimization variable is a single parameter (panel -c-) initialized at 1, which defines the diagonal Mahalanobis matrix for the reward: $\mathrm{u}>0$ s.t. $\mathrm{U}=\mathrm{u}\cdot\mathrm{I}_{n}$. Panel -a- illustrates the evolution of the outer loss (NLL evaluated on the optimal policy from the inner PG loop), showing clear improvement relative to the initialization at 1 (which corresponds to the Euclidean distance). Additionally, the optimization parameter $\mathrm{u}$ converges to a value close to the theoretical optimum $\mathrm{u}^{\star}=\frac{\lambda n}{2\operatorname{Tr}(\Sigma)}$, as derived in 4.2. This convergence is further supported by panel -e-, which plots the outer loss landscape as a function of the reward parameter $\mathrm{u}$, revealing a roughly convex landscape with a global minimum near the theoretical optimum. These results validate our intuition from the tractable case discussed in Section 4, even in the more general setting of MLP-based policies and stochastic optimization solvers within the implicit differentiation framework.

### v3

# From Data to Rewards: a Bilevel Optimization Perspective on Maximum Likelihood Estimation

@@ 4.1 Theoretical analysis
[Closed-form solution] Under assumption 4.1 and 4.2, the Bi-O problem
@@
The proof of assumption 4.2 is deferred to Section A.1.
@@
Under the assumptions of assumption 4.2, the optimal parameters $\theta^{\star}_{\mathrm{U}^{\star}}$ obtained from the lower-level problem with $\mathrm{U}^{\star}=\frac{\lambda}{2}\Sigma^{-1}$ minimize the reverse KL divergence between $\hat{p}_{\theta}$ and $q$:
@@ 4.2 Empirical validation
In this section, we empirically evaluate the theoretical results from Section 4.1. To this end, we generate synthetic data that satisfy assumption 4.1: $\mathcal{D}=\{({\bf x}_{i},{\bf y}_{i})\}_{i=0}^{N}$, where ${\bf x}_{i}\sim\mathcal{U}([-5,5]^{n})$ ($\mathcal{U}$ denotes the uniform distribution), and ${\bf y}_{i}\sim q(.|{\bf x}_{i}):=\mathcal{N}(\Lambda{\bf x}_{i},\Sigma)$ with diagonal covariance matrix $\Sigma=\beta^{2}\mathrm{I}_{n}$ and $\beta>0$. For the model $\hat{p}_{\theta}$, we relax the linearity and homoscedasticity assumptions by considering a neural network that parametrizes a Gaussian distribution, in which both the mean function and the diagonal covariance matrix depend on the input: $\hat{p}_{\theta}(.|{\bf x}_{i}):=\mathcal{N}(\mu_{\theta}({\bf x}_{i}),\sigma^{2}_{\theta}({\bf x}_{i})\mathrm{I}_{n})$. We compare baselines trained with the negative log-likelihood (NLL) and mean squared error (MSE) losses against PG variants, using either a negative squared distance reward $\mathrm{U}=\mathrm{I}_{n}$ or the optimal reward function derived in assumption 4.2 with $\mathrm{U}^{\star}=\frac{\lambda n}{2\operatorname{Tr}(\Sigma)}\mathrm{I}_{n}$.
@@
Figure 2: Learned distributions comparison on a single data point. The PG loss paired with the optimal reward function in assumption 4.2 shows optimal convergence, even when compared with the baseline directly optimizing the NLL.
@@ 5.2 Empirical validation
Fig. 3 presents the results of running an implicit differentiation solver for 100 outer iterations, each with 50 inner iterations, and a learning rate of $10^{-2}$ for both optimization loops. The outer optimization variable is a single parameter (panel -c-) initialized at 1, which defines the diagonal Mahalanobis matrix for the reward: $\mathrm{u}>0$ s.t. $\mathrm{U}=\mathrm{u}\cdot\mathrm{I}_{n}$. Panel -a- illustrates the evolution of the outer loss (NLL evaluated on the optimal policy from the inner PG loop), showing clear improvement relative to the initialization at 1 (which corresponds to the Euclidean distance). Additionally, the optimization parameter $\mathrm{u}$ converges to a value close to the theoretical optimum $\mathrm{u}^{\star}=\frac{\lambda n}{2\operatorname{Tr}(\Sigma)}$, as derived in assumption 4.2. This convergence is further supported by panel -e-, which plots the outer loss landscape as a function of the reward parameter $\mathrm{u}$, revealing a roughly convex landscape with a global minimum near the theoretical optimum. These results validate our intuition from the tractable case discussed in Section 4, even in the more general setting of MLP-based policies and stochastic optimization solvers within the implicit differentiation framework.